OpenStudy (anonymous):

2x + y = 8 x + y = 4 The lines whose equations are given intersect at (4, 0) (0, 4) all points on the line

OpenStudy (anonymous):

\[2x + y = 8\]\[ x + y = 4 \] subtract the second equation from the first in a column you get \[x=4\]

OpenStudy (anonymous):

since \(x=4\) you know \(4+y=4\) and so \(y=0\) your answer is \((4,0)\)